OpenStudy (anonymous):

I really need help, please! I am having the worst time with these problems. I can't remember how to do them and I have been trying to work on this work sheet for quite a while now, which is putting me behind in my class. I could really use some help! I just need someone to explain how to solve it and then I think I can get the rest. I have 11 problems on my work sheet. I have posted the question below. Thank you! :)

OpenStudy (anonymous):

Sure ! But say that please !:)

OpenStudy (texaschic101):

where is the question ?

OpenStudy (anonymous):

It's right there. It took a minute to upload. And E.ali, say what? lol Thanks guys! :)

OpenStudy (texaschic101):

1. your two points are (-5,5) and (1,-2) do you agree ?

OpenStudy (anonymous):

I'd say so, yes.

OpenStudy (anonymous):

K ! To find the distance between two point we use this formula : \[\sqrt{(x1-x2)^2+(y1-y2)^2}\]

OpenStudy (anonymous):

Fisrt find the place of the points and write them please :)

OpenStudy (anonymous):

@Madds96

OpenStudy (texaschic101):

the square root of (x2 - x1)^2 + (y2 - y1)^2 (-5,5) x1 = -5 and y1 = 5 (1,-2) x2 = 1 and y2 = -2 lets plug it into the formula square root of (1 - 5)^2 + (-2 - 5)^2 square root of (-4)^2 + (-7)^2 square root of 16 + 49 square root of 65 = 8.06 rounded to 8.1

OpenStudy (texaschic101):

do the same thing for the second problem

OpenStudy (anonymous):

I'm honestly really confused still...lol I'm sorry. I'm awful at math... :(

OpenStudy (texaschic101):

square roots and powers can be confusing.

OpenStudy (texaschic101):

the points on the second graph are what ?

OpenStudy (anonymous):

CAn you find the Coordinate ?

OpenStudy (anonymous):

(-5, 1), (4, 0) I believe. :)

OpenStudy (texaschic101):

correct.. lets put them in the formula (-5,1) x1 = -5 and y1 = 1 (4,0) x2 = 4 and y2 = 0 sq rt (x2 - x1)^2 + (y2 - y1)^2 sq rt (4 -(-5)^2 + (0 - 1)^2 can you finish ?

OpenStudy (anonymous):

Is this one sq rt (x2 - x1)^2 + (y2 - y1)^2 1.41?

OpenStudy (texaschic101):

sq rt (4 + 5)^2 + (0 - 1)^2 sq rt 9^2 + -1^2 sq rt 81 + 1 sq rt 82 = 9.05 rounds to 9.1 do you understand how I got this ?

OpenStudy (texaschic101):

sq rt (x2 - x1)^2 + (y2 - y1)^2 is the distance formula

OpenStudy (anonymous):

I understand how you got this part; sq rt (4 + 5)^2 + (0 - 1)^2 sq rt 9^2 + -1^2 sq rt 81 + 1 sq rt 82 = 9.05 rounds to 9.1 But I don't understand how you got that out of this part; (-5,1) x1 = -5 and y1 = 1 (4,0) x2 = 4 and y2 = 0 sq rt (x2 - x1)^2 + (y2 - y1)^2 sq rt (4 -(-5)^2 + (0 - 1)^2

OpenStudy (texaschic101):

First find your points. Then you have to label your points. (x1, y1) and (x2,y2) you still with me ?

OpenStudy (texaschic101):

your points are (-5,1) -- this is x1 and y1 (4,0) -- this is x2 and y2 understand ?

OpenStudy (anonymous):

I think so. so it would be: (x[-5], y[1]), (x[4], y[0])?

OpenStudy (texaschic101):

yes...you have 2 sets of points. First set contains x1 and y1 and 2nd set contains x2 and y2

OpenStudy (anonymous):

Ok. I understand that part. :)

OpenStudy (texaschic101):

(-5,1) can either be set 1 or set 2, it doesn't matter. Just be sure and label each point correctly

OpenStudy (anonymous):

Ok. I'll make it be set 1, since it comes first I'll be able to understand it better.

OpenStudy (texaschic101):

whatever makes it easier for you

OpenStudy (texaschic101):

do you have any more questions ?

OpenStudy (anonymous):

Ok, for (-5, 1) would it be written like this; (-5, 1) * 1 = -5?

OpenStudy (texaschic101):

no....where did that come from ?

OpenStudy (texaschic101):

why are you multiplying it by 1 ?

OpenStudy (anonymous):

Oh wait, I read one of the things above wrong... :P

OpenStudy (texaschic101):

do you understand about labeling the points and subbing it into the formula ?

OpenStudy (texaschic101):

if not, this may help http://www.mathwarehouse.com/algebra/distance_formula/index.php

OpenStudy (texaschic101):

good luck :)

OpenStudy (anonymous):

I'm not sure. This is and example that my teacher gave me of how I have to write it out. (-6, -10), (-2, -10) d = \[\sqrt{(x2 - x1)^2 + (y2 - y1)^2}\] \[\sqrt{(-2 - -6)^2 + (-10 - -10)^2}\] \[\sqrt{(-2 + 6)^2 + (-10 + 10)^2}\] \[\sqrt{4^2 + 0^2}\] \[\sqrt{16 = 4}\]

OpenStudy (anonymous):

Realyyyyyyyyyyyyyyyy right !:)

OpenStudy (anonymous):

The problem is, it all looks like gibberish to me.

OpenStudy (texaschic101):

that is correct

OpenStudy (anonymous):

WAIT!!! I think I might have it!

OpenStudy (anonymous):

Would a graph with the plots (5, -5), (1, -2) be started like this?: \[\sqrt{(1 - 5)^2 + (-2 - -5)^2}\]\[\sqrt{(1 + 5) ^2 + (-2 + -5)^2}\]

OpenStudy (anonymous):

\[\sqrt{(6)^2 + (-7)^2}\]

OpenStudy (texaschic101):

not quite sq rt (1 - 5)^2 + (-2 - (-5)^2 sq rt (1 - 5)^2 + (-2 + 5)^2 sq rt -4^2 + 3^2

OpenStudy (anonymous):

Oh, so if its 2 negatives, you change the subtraction to addition, but if its 2 positive's, you leave the subtraction?

OpenStudy (texaschic101):

do not distribute anything through the parenthesis

OpenStudy (texaschic101):

if it is 2 negatives multiplied together, change it to a positive. If it is 2 positives, you do not change anything

OpenStudy (texaschic101):

so yes.....you leave the subtraction

OpenStudy (anonymous):

So would the final answer to the above problem be: sq rt 16 + 9

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

Ok. I'm pretty sure that I get how to do it now. Can we do another one to be sure?

OpenStudy (texaschic101):

of course

OpenStudy (anonymous):

Ok, lets do #2. I believe that the points are (-5, 1), (0, 4)

OpenStudy (texaschic101):

(-5,1) and (4,0)

OpenStudy (anonymous):

sq rt (0 - -5)^2 + (4 - 1)^2, right?

OpenStudy (texaschic101):

wait...your right (4,0)

OpenStudy (texaschic101):

correct

OpenStudy (texaschic101):

wait......mistake......check your points.............it is (4,0)

OpenStudy (anonymous):

Oh my bad... so sq rt (4 - -5) + (0 - 1)?

OpenStudy (texaschic101):

yes....sorry about that

OpenStudy (anonymous):

It's not your fault, it's mine. :)

OpenStudy (anonymous):

So then, (4- -5)^2 + (0 - 1)^2?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

sq rt 81 + 1?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

And its done now?

OpenStudy (texaschic101):

no.....sq rt 82 = 9.05 which rounded to the nearest tenths = 9.1

OpenStudy (anonymous):

Opps, I forgot about that step... So to do it, I add the 81 & 1 = 82. Then find the sq rt. 9.05, round to 9.1?

OpenStudy (texaschic101):

this helps... http://www.math.com/students/calculators/source/square-root.htm

OpenStudy (texaschic101):

do not find the square root of 9.05.....that is the answer, but you have to round the answer to the tenths place, which is 9.1

OpenStudy (texaschic101):

find the square root of 82

OpenStudy (anonymous):

Thats what I meant in my head, I typed it wrong. :P

OpenStudy (texaschic101):

no problem.....if I had a dollar for every mistake I made, I would be rich

OpenStudy (anonymous):

Haha so would I! :) Thank you for all of the help, I really appriciate it!

OpenStudy (texaschic101):

no problem...thats what I am here for

OpenStudy (texaschic101):

have a good day

OpenStudy (anonymous):

Thanks, you too!